OpenStudy (anonymous):

solve: 3^(2x +1) = 70

OpenStudy (anonymous):

let me write it out for you.. give me amin

OpenStudy (anonymous):

\[3^{2x+1} = 70\] \[\log 3^{2x+1} = \log 70\] \[(2x+1) \log3 = \log70\] \[2x = (\log70/\log3) -1\] x =( log70/log3 - 1) /2

OpenStudy (anonymous):

then just put that into ur calculator and voila :)

OpenStudy (anonymous):

app 1.43

OpenStudy (anonymous):

is there a way it can be done without logs?